Denise bought a box with both chocolate and maple donuts. There were 3 fewer maple donuts than chocolate donuts, and there were 13 donuts in all. How many chocolate donuts did Denise buy?
1
Expert's answer
2020-02-06T14:03:03-0500

As we want to know how many chocolate donuts did Denise buy.

Let X = chocolate donuts that Denise bought. Then maple donuts is X-3 as there were 3 fewer maple donuts than chocolate donuts. So in total we have chocolate + maple donuts or X + (X-3) and it exactly equals 13 according to the task.

The equality we've got:

X + (X-3) = 13

Open the brackets (as there the plus before the bracket it doesn't change signs of the expression in brackets)

X + X-3 = 13

After adding and carrying -3 over equal sign we get

2*X = 16

Dividing both sides by 2

So the X = 16 : 2 = 8

The answer is Denise bought 8 chocolate donuts.